How much pure dye must be added to 4 gal of a 25% dye solution to increase the solution to 40% ?(hint:Pure Dye is 100% dye.)
Answer by stanbon(75887)

How much pure dye must be added to 4 gal of a 25% dye solution to increase the solution to 40% ?
----------------------
Equation:
dye + dye = dye
0.25*4 + x = 0.40(4+x)
1 + x = 1.6 + 0.4x
0.6 = 0.6x
x = 1 gal (amt. of 100% dye that must be added)
=========================================================
Cheers,
Stan H.